Mrs. Brady invests a total of $8218 in two savings accounts. One account yields 8.5% simple interest and the other 9% simple interest. She earned a total of $706.47 interest for the year. How much was invested in each account?
Answer by lwsshak3(11628)

Mrs. Brady invests a total of $8218 in two savings accounts. One account yields 8.5% simple interest and the other 9% simple interest. She earned a total of $706.47 interest for the year. How much was invested in each account?
***
let x= amt invested in 8.5% account
8218-x= amt invested in 9.0% account
.08x+.09(8218-x)=706.47
.08x+739.62-.09x=706.47
.01x=33.15
x=3315
8218-x=4903
amt invested in 8.5% account=$3315
amt invested in 9.0% account=$4903